dejamarie:

The local skating rink pays Mary a fixed rate per pupil plus a base amount to work as a skating instructor. She earns $90 for instructing 15 students on Monday afternoon. Last Friday, she earned $62 for working with 8 students. Lisa is also a skating instructor. She receives half the base amount that Mary does, but she is paid twice as much per student. Who would earn more money instructing a class of 20 students? Lisa would earn $120 more than Mary. Lisa would earn $65 more than Mary. Mary would earn $15 more than Lisa. Mary would earn $10 more than Lisa.

jhonyy9:

any idea ?

justus:

y = 4x + 30 y = 4(20) + 30 = 110 y = 8x + 15 y =8(20) + 15 = 175

justus:

The difference between both Mary and Lisa's salary = 175-110=65 dollars.

justus:

Now you can look In the answer choices and see which answer suits best. Good luck.